# MAPF-GPT: Imitation Learning for Multi-Agent Pathfinding at Scale

### v3

@@ Abstract
Multi-agent pathfinding (MAPF) is a challenging computational problem that typically requires to find collision-free paths for multiple agents in a shared environment. Solving MAPF optimally is NP-hard, yet efficient solutions are critical for numerous applications, including automated warehouses and transportation systems. Recently, learning-based approaches to MAPF have gained attention, particularly those leveraging deep reinforcement learning. Following current trends in machine learning, we have created a foundation model for the MAPF problems called MAPF-GPT. Using imitation learning, we have trained a policy on a set of pre-collected sub-optimal expert trajectories that can generate actions in conditions of partial observability without additional heuristics, reward functions, or communication with other agents. The resulting MAPF-GPT model demonstrates zero-shot learning abilities when solving the MAPF problem instances that were not present in the training dataset. We show that MAPF-GPT notably outperforms the current best-performing learnable-MAPF solvers on a diverse range of problem instances and is efficient in terms of computation (in the inference mode).
Codehttps://github.com/CognitiveAISystems/MAPF-GPT
@@ Introduction
Multi-agent pathfinding (MAPF) (Stern et al. 2019) is a combinatorial computational problem that asks to find a set of paths for the agents that operate in a shared environment, such that accurately following these paths does not lead to collisions and, preferably, each agent reaches its specified goal as soon as possible. On the one hand, even under simplified assumptions, such as graph representation of the workspace, discretized time, and uniform duration of actions, optimally solving MAPF is NP-Hard (Surynek 2010). On the other hand, efficient MAPF solutions are highly demanded in numerous real-world applications, such as automated warehouses (Li et al. 2021), railway scheduling (Svancara and Barták 2022), transportation systems (Mironov et al. 2023), etc. This has resulted in a noticeable surge of interest in MAPF and the emergence of a large body of works devoted to this topic.
@@
Recently, learning-based MAPF solvers have come on stage (Skrynnik et al. 2021; Alkazzi and Okumura 2024). They mostly rely on deep reinforcement learning and typically involve additional components to enhance their performance, such as single-agent planning, inter-agent communication, etc. Meanwhile, in the realm of machine learning, currently, the most impressive progress is driven by self-supervised learning (at scale) on expert data and employing transformer-based architectures (Vaswani et al. 2023). It is this combination that recently led to the creation of the seminal large-language models (LLMs) and (large) vision-language models (VLMs) that achieve an unprecedented level of performance in text generation (dialogue) systems (Dubey et al. 2024), image and video generation (Liu et al. 2024; Zhu et al. 2024), etc. Moreover, such a data-driven approach has become widespread in robotics and reinforcement learning, where an agent’s imitation policy is trained based on a variety of the expert trajectories (Chen et al. 2021; Chi et al. 2023). Thus, in this work, we are motivated by the following question: Is it possible to create a strong learnable MAPF solver (that outperforms state-of-the-art competitors) purely on the basis of supervised learning on expert data omitting additional decision-aiding routines like communication, single-agent planning, etc.? Our answer is positive.
@@
To create our learning-based MAPF solver, which we name MAPF-GPT, we first design a vocabulary of terms, called tokens in machine learning, that is used to describe any observation an individual agent may perceive and any action it can perform while navigating in the environment. Next, we create a large and diverse dataset of expert data, i.e., successful sub-optimal MAPF solutions, utilizing state-of-the-art MAPF solver. Consequently, we convert these MAPF solutions into the sequences of (observation, action) tuples, encoded with our tokens, and utilize a transformer-based non-autoregressive neural network to learn to predict the correct action provided with the observation. Since sub-optimal expert actions are present in the dataset, using a simple cross-entropy loss function allows the agent to update the parameters of the transformer-based policy. In our extensive empirical evaluation, we show that MAPF-GPT notably overpasses the current best-performing learnable-MAPF solvers (without any usage of additional planning or communication mechanisms), especially when it comes to out-of-distribution evaluation, i.e., evaluating the solvers on the problem instances that are not similar to the ones used for training (a common bottleneck for learning-based solvers). We also report ablation studies and evaluate MAPF-GPT in the other type of MAPF, i.e. the Lifelong MAPF (both in zero-shot and finetuning regimes).
@@
  • We present the largest MAPF dataset for decision-making, consisting of 1 billion (observation, action) pairs.
@@
  • We extensively study MAPF-GPT and compare it with state-of-the-art decentralized learning-based approaches, showing MAPF-GPT’s superior performance across a wide range of tasks, including out-of-distribution ones, along with better runtime efficiency.
@@
Multiagent pathfinding
Several orthogonal approaches to tackle MAPF can be distinguished. First, dedicated rule-based MAPF solvers exist that are tailored to obtaining MAPF solutions fast, yet no bounds on the suboptimality of the resultant solutions are guaranteed (Okumura 2023; Li et al. 2022). Second, reduction-based approaches to obtain optimal MAPF solutions are widespread. They convert MAPF to some other well-established computer science problems, e.g., minimum-flow on graphs, boolean satisfiability (SAT), and employ off-the-shelve solvers to obtain the solution of this (converted) problem (Yu and LaValle 2013; Surynek et al. 2016). Next, a plethora of the search-based MAPF solvers exist (Sharon et al. 2015, 2013; Wagner and Choset 2011). They explicitly rely on the graph-search techniques to obtain MAPF solutions and often may provide certain desirable guarantees, e.g., guarantees on obtaining optimal or bounded suboptimal solutions (meanwhile, simplistic search-based planners that lack strong guarantees, like prioritized planning (Ma et al. 2019), are also widespread).
There were also multiple attempts to apply machine and reinforcement learning to solve MAPF. One of the first such successful solvers was PRIMAL (Damani et al. 2021) that demonstrated how MAPF problem can be solved in a decentralized manner utilizing machine learning. The recent learnable MAPF solvers such as SCRIMP (Wang et al. 2023), DCC (Ma, Luo, and Pan 2021), Follower (Skrynnik et al. 2024b), to name a few, typically rely on reinforcement learning and often rely on additional modules, like the communication one, to solve the problem at hand. Orthogonally to these approaches we rely purely on imitation learning from (a large volume) of expert data.
@@
Offline deep reinforcement learning develops a policy based on previously collected data without interacting with the environment during the training phase (Levine et al. 2020). This approach allows for the utilization of large amounts of pre-collected data to develop a robust policy. There are numerous effective offline RL approaches, such as CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2022), and TD3+BC (Fujimoto and Gu 2021). Modern approaches often involve using transformers as the architectural backbone. One popular approach is the Decision Transformer (DT), which models the behavior of an expert by conditioning on the desired outcomes, thereby integrating reward guidance directly into the decision-making process. In multi-agent scenarios, there is less diversity in offline RL methods; however, a multi-agent adaptation of the DT exists, known as MADT (Meng et al. 2021).
The application of offline RL to solving challenging tasks often requires additional techniques. One commonly used technique is behavioral cloning, which involves supervised learning on large datasets. A notable example of this approach is AlphaStar (Vinyals et al. 2019), where the authors pre-trained a large model (including transformer blocks) on expert gameplay data from StarCraft II and then fine-tuned it using RL. Similarly, VPT (Video Pretraining) (Baker et al. 2022) addresses the task of playing Minecraft by proposing a foundation model trained on a large video dataset collected from YouTube. In the case of AlphaGo (Silver et al. 2016), pretraining on human gameplay data was used before the development of AlphaZero (Silver et al. 2017). Additionally, transformers trained with a combination of behavioral cloning and offline RL losses have demonstrated grandmaster-level chess performance without the use of search techniques (Ruoss et al. 2024).
Multiagent imitation learning (MAIL)
@@
Imitation learning and learning from demonstration are actively used in multi-agent systems (Tang et al. 2024; Liu and Zhu 2024). MAIL refers to the problem of agents learning to perform a task in a multi-agent system through observing and imitating expert demonstrations without any knowledge of a reward function from the environment. It has gained particular popularity in the tasks of controlling urban traffic and traffic lights at intersections (Bhattacharyya et al. 2018; Huang et al. 2023) due to the presence of a large amount of data collected in real conditions and a high-quality simulator (such as Sumo (Lopez et al. 2018)). Among the methods in the field of MAIL, it is possible to note works using the Bayesian approach (Yang et al. 2020), generative adversarial methods (Song et al. 2018; Li et al. 2024), statistical tools for capturing multi-agent dependences (Wang et al. 2021), low-rank subspaces (Shih, Ermon, and Sadigh 2022), latent multi-agent coordination models (Le et al. 2017), decision transformers (Meng et al. 2021), etc. Demonstrations are also often used as an element of pre-training in game tasks (for example, pre-training the value function for AlphaGo (Silver et al. 2016)) and in MAPF tasks (as in SCRIMP (Wang et al. 2023)). However, despite the listed works in this area, a single foundation model has not yet been proposed, the imitation learning of which already gives high results in multi-agent tasks and does not require an additional stage of online learning in the environment. This is largely due to the complexity of behavioral multi-agent policies in various tasks (such as in Starcraft (Samvelyan et al. 2019) and traffic control) and the lack of large datasets of expert trajectories, which are necessary for effective training of foundation models. In this regard, the MAPF task is a convenient testbed for investigating transformer foundation models in a multi-agent setting, which will provide additional insights for creating such models in other applications, and our work also provides a large dataset for training MAPF models.
@@
The classical variant of the MAPF problem is defined by a tuple $(n,\mathcal{G=(V,E)},S=\{s_{1},...,s_{n}|~{}s_{i}\in\mathcal{V}\},G=\{g_{1},...,g_{n}|~{}g_{i}\in\mathcal{V}\})$, where $n$ is the number of agents acting in the shared workspace which is represented as an undirected graph $\mathcal{G}$. At each time step, an agent is assumed to either move from one vertex to the other or wait at the current vertex (the duration of both actions is uniform and equals $1$ time step). The plan for the $i$-th agent, $pl_{i}$, is a sequence of moves, s.t., each move starts where the previous one ends. Two distinct plans are said to contain a vertex/edge conflict if a time step exists, s.t., the agents following these plans occupy the same graph vertex/traverse the same edge in opposite directions.
The problem is to find a set of $n$ plans, $Pl=\{pl_{1},...,pl_{n}\}$, s.t. each $pl_{i}$ starts at $s_{i}$, ends at $g_{i}$ and each pair of plans in $Pl$ is conflict-free. The objective to be minimized is typically defined as $SoC(Pl)=\sum_{i=1}^{n}cost(pl_{i})$ (called the Sum-Of-Costs) or as $MS(Pl)=\max_{i=1,...,n}cost(pl_{i})$ (called the Makespan), where $cost(pl_{i})$ is the cost of the individual plan which equals the time step when the agent reaches its goal vertex (and does not move away further on).
@@
Despite MAPF is typically considered to be a planning problem as defined above, it can also be considered as a sequential decision-making (SDM) problem. Within the SDM framework, the problem is to construct a policy, $\pi$, that is a function that maps the current state (current positions of all agents in the graph) to a (joint) action $\mathbf{a}=a_{i}\times...\times a_{n}$, where $a_{i}\in A_{i}$ and $A_{i}$ is the set of possible actions for agent $i$. When $\pi$ is obtained, it is invoked sequentially until either all agents reach their goals or the threshold on the number of time steps, $t_{max}$, is reached.
For better scalability, the decision-making policy might be decentralized, i.e., each agent chooses its action independently of the other agents. In practice, decentralized agents typically don’t have access to the global state of the environment, i.e., positions of the other agents, but rather rely on local observation, $o_{t}$. For example, if the underlying graph is a 4-connected grid, then the local observation may be a $(2r+1)\times(2r+1)$ patch of the grid centered at the agent’s current position (where $r$ is the observation radius) and the latter observes only the agents that are within this patch. A sequence of individual observations and actions forms the agent’s history: $h_{t}=\{o_{1},a_{1},o_{2},a_{2},...,o_{t-1},a_{t-1},o_{t}\}$, where $o_{k}$ and $a_{k}$ denote action and observation at time step $k$. This history is typically used to reconstruct a markovian state of the environment via some approximator $f$: $s_{t}\approx f(h_{t})$ (e.g. $f$ can be represented as a neural network).
@@
To construct (learn) a decision making policy $\pi$ imitation learning relies on an expert policy $\pi^{b}$, which is used to collect a set of trajectories: $\mathcal{D}=\{traj\}$, where each trajectory is composed of the observations and actions: $traj=\{o^{b}_{1},a^{b}_{1},\dots,o^{b}_{L},a^{b}_{L}\}$. Intuitively, in MAPF context, $\mathcal{D}$ represents the expert knowledge on how an agent should behave under different circumstances (it can be obtained running a well-established MAPF solver on a range of problem instances).
@@
where $a_{j}\sim\pi_{\theta}(s_{j})$, $a^{b}_{j}\sim\pi(s^{b}_{j})$, and $\mathcal{L}$ is a loss function, which is selected depending on the action space. In the considered case, when the action space is discrete, cross-entropy is widely used as $\mathcal{L}$.
@@
Figure 1: The general pipeline of the MAPF-GPT: (1) Creating MAPF scenarios. (2) Generating ground truth data, i.e. MAPF solutions using an expert solver. (3) Transforming the solutions to the (observation, action) pairs and tokenization of the observations, which converts them into a format suitable for transformer architectures. (4) Executing the training loop, where observation/action pairs are sampled from the dataset, and the model is trained using cross-entropy loss.
@@
A large, curated dataset is crucial for any data-driven method including ours. To create the set of training instances we used POGEMA (Skrynnik et al. 2024a), a versatile tool for developing learnable MAPF solvers that includes utilities to generate maze-like maps and maps with random obstacles, as well as to create MAPF instances from them (i.e. assigning start-goal locations). For our purposes we generated 10K of maze-like maps and 2.5K random maps and further created 3.75M different problem instances on these maps. The size of the maps varied from $17\times 17$ to $21\times 21$, the number of the agents was $16$, $24$, or $32$. Please note that as we aim to create an individual policy to solve MAPF in a decentralized fashion (i.e. each agent makes its own decision on how to move based on its local observation) it is not the size of the maps that actually matters but rather the density, i.e. the ratio of the free space to the space occupied by the agents. We use moderate and considerably high densities to make the agents face challenging patterns requiring coordination (especially on the maze-like maps).
@@
To create expert data we used a recent variant of LaCAM (Okumura 2024, 2023), a state-of-the-art MAPF solver that is tailored to quickly find a solution and iteratively enhance it while having a time budget. As we need to solve a large number of MAPF instances (i.e. 3.75M) we set this time budget to be $10$ seconds.
The output of LaCAM on a single MAPF instance is the set of the individual plans. We, further, trace each plan and reconstruct an agent’s (local) observations in order to form (observation, action) pairs. We use the following post-processing to filter out some of them. First, if several pairs share the exactly the same observation we keep only one of them (picked randomly). Second, we observe that the fraction of the pairs when an agents waits at the goal is very high, because in many cases numerous agents wait for long times for other agents to reach their goals and just stand still. In fact, almost $40\%$ of the actions in the expert data were the waiting ones. To remove this imbalance we discard $80\%$ of wait-at-target actions. We end up with 900M of (observation, action) pairs from the maze-like maps, and 100M – from the random ones. A 9:1 proportion was chosen due to the maze maps possessing more challenging layouts with numerous narrow passages that require a high degree of cooperation between the agents.
We believe that the obtained dataset composed of 1B (observation, action) pairs is currently the largest dataset of such kind and may bring value to the other researchers developing learnable MAPF solvers. Additional technical details are presented in the Appendix.
@@
Figure 2: Tokenization process for the decentralized MAPF model. The input consists of 256 tokens, but the example shown uses fewer tokens for clarity and visibility.
Tokenization can be thought of as the process of transforming the data, (observation, action) pairs in our case, into the sequence of special symbols, tokens, to be further fed to the neural network that is trained to predict a single token, i.e. action, from the sequence of the input tokens, i.e. the ones that encode the observation. The input tokens are typically referred to as the context. We now wish describe how our context, i.e. the observation, is structured.
@@
The second part of the observation contains data about the agent itself and the nearby ones. The information about each agent consists of the coordinates of its current and goal locations, actions history, i.e., the actions that were made in the previous $k$ steps, and the action that is preferable w.r.t. the agent’s individual cost-to-go map – the so-called greedy action. Please note that there may be cases where more than one action leads to a decrease in cost-to-go. Thus, we use special markers to indicate these ”multi-direction” greedy actions (e.g., ”up-right”).
@@
The rest of the input ($135$ tokens) is devoted to the information about agents. As it’s important to consider only the agents that can potentially influence the egocentric agent, we consider only the ones that are located in the $11\times 1$ 11 field of view at the current time step. The information about each agent is encoded via $10$ tokens: $2$ for the current position, $2$ for goal location, $5$ for action history, and $1$ for the next greedy action. Thus, we are able to encode the information about at most $13$ agents, including the egocentric one. The rest of the $5$ tokens in the input are encoded with the empty token. In case there are not enough agents in the local field of view, the information about missing agents is also filled with empty tokens. The information about agents is sorted based on the distance to the egocentric agent, i.e., the information about the egocentric agent itself always goes first, as its distance is always 00.
@@
Figure 3: Performance of MAPF approaches on Random, Mazes, Warehouse and MovingAI maps, based on Success Rate (higher is better). The shaded area indicates $95\%$ confidence intervals.
Figure 4: Performance of MAPF approaches on Random, Mazes, Warehouse and MovingAI maps, based on SoC metric relative to LaCAM performance (lower is better).
@@
As the model backbone, we used a modern decoder-only transformer (Brown et al. 2020). We used a softmax layer to parameterize the discrete probability distribution. To sample an action, we then used multinomial sampling. The length of the input sequence (context size) of the model is $256$. The output size of the model is $5$ since each agent has $5$ available discrete actions. We used learnable position embeddings. Our largest model has $85$M parameters. Besides it, we also trained models with $6$M and $2$M parameters. We don’t use causal masking, which is common practice in NLP (Radford et al. 2019), since the model predicts only a single action ahead in a non-autoregressive manner. To speed up training, we used the flash attention technique (Dao 2024).

### v5

@@ Abstract
Multi-agent pathfinding (MAPF) is a problem that generally requires finding collision-free paths for multiple agents in a shared environment. Solving MAPF optimally, even under restrictive assumptions, is NP-hard, yet efficient solutions for this problem are critical for numerous applications, such as automated warehouses and transportation systems. Recently, learning-based approaches to MAPF have gained attention, particularly those leveraging deep reinforcement learning. Typically, such learning-based MAPF solvers are augmented with additional components like single-agent planning or communication. Orthogonally, in this work we rely solely on imitation learning that leverages a large dataset of expert MAPF solutions and transformer-based neural network to create a foundation model for MAPF called MAPF-GPT. The latter is capable of generating actions without additional heuristics or communication. MAPF-GPT demonstrates zero-shot learning abilities when solving the MAPF problems that are not present in the training dataset. We show that MAPF-GPT is able to outperform the state-of-the-art learnable MAPF solvers on a diverse range of problem instances and is computationally efficient during inference.
Project Pagehttps://sites.google.com/view/mapf-gpt/
@@ Introduction
Multi-agent pathfinding (MAPF) (Stern et al. 2019) is a combinatorial computational problem that asks to find a set of paths for the agents that operate in a shared environment, such that accurately following these paths does not lead to collisions and, preferably, each agent reaches its specified goal as soon as possible. On the one hand, even under simplified assumptions, such as a graph representation of the workspace, discretized time, uniform duration of actions, optimally solving MAPF is NP-Hard (Surynek 2010). On the other hand, efficient MAPF solutions are highly demanded in numerous real-world applications, such as automated warehouses (Li et al. 2021), railway scheduling (Svancara and Barták 2022), transportation systems (Li et al. 2023), etc. This has resulted in a noticeable surge of interest in MAPF and the emergence of a large body of research devoted to this topic.
@@
Recently, learning-based MAPF solvers have come on stage (Skrynnik et al. 2021; Alkazzi and Okumura 2024; Skrynnik et al. 2023). They mostly rely on deep reinforcement learning and typically involve additional components to enhance their performance, such as single-agent planning, inter-agent communication, etc. Meanwhile, in the realm of machine learning, currently, the most impressive progress is driven by self-supervised learning (at scale) on expert data and employing transformer-based architectures (Vaswani et al. 2023). It is this combination that recently led to the creation of the seminal large-language models (LLMs) and (large) vision-language models (VLMs) that achieve an unprecedented level of performance in text, image and video generation (Dubey et al. 2024; Liu et al. 2024; Zhu et al. 2024). Moreover, such data-driven approach has become widespread in robotics, where an imitation policy is trained based on a variety of expert trajectories (Chen et al. 2021; Chi et al. 2023). Thus, in this work we are motivated by the following question: Is it possible to create a strong learnable MAPF solver (that outperforms state-of-the-art competitors) purely on the basis of supervised learning (at scale) on expert data omitting additional decision-aiding routines? Our answer is positive.
@@
To create our learning-based MAPF solver, which we name MAPF-GPT, we, first, design a vocabulary of terms, called tokens in machine learning, that are used to describe any observation an individual agent may perceive and any action it may perform. Next, we create a diverse dataset of expert data, i.e., successful MAPF solutions, utilizing a state-of-the-art MAPF solver. Consequently, we convert these MAPF solutions into sequences of observation-action tuples, encoded with our tokens, and utilize a transformer-based non-autoregressive neural network to learn to predict the correct action provided with the observation. In our extensive empirical evaluation, we show that MAPF-GPT notably overpasses the current best-performing learnable-MAPF solvers (without any usage of additional planning or communication mechanisms), especially when it comes to out-of-distribution evaluation, i.e., evaluating the solvers on the problem instances that are not similar to the ones used for training (a common bottleneck for learning-based solvers). We also report ablation studies and evaluate MAPF-GPT in another type of MAPF, i.e. the Lifelong MAPF (both in zero-shot and fine-tuning regimes).
@@
  • We present the largest MAPF dataset for decision-making, containing 1 billion observation-action pairs.
@@
  • We extensively study MAPF-GPT and compare it with state-of-the-art decentralized learning-based approaches, showing MAPF-GPT’s high performance across a wide range of tasks, along with better runtime efficiency.
@@
Multi-agent pathfinding
Several orthogonal approaches to tackle MAPF can be distinguished. First, dedicated rule-based MAPF solvers exist that are tailored to obtaining MAPF solutions fast, yet no bounds on their costs are guaranteed (Okumura 2023; Li et al. 2022). Second, reduction-based approaches to obtain optimal MAPF solutions are widespread. They convert MAPF to some other well-established computer science problem, e.g., minimum-flow on graphs, boolean satisfiability (SAT), and employ an off-the-shelf solver to obtain the solution of this problem (Yu and LaValle 2013; Surynek et al. 2016). Next, a plethora of search-based MAPF solvers exist (Sharon et al. 2015, 2013; Wagner and Choset 2011). They explicitly rely on graph-search techniques to obtain MAPF solutions and often may provide certain desirable guarantees, e.g., optimal or bounded suboptimal solutions. Meanwhile, simplistic search-based planners that lack strong guarantees, like prioritized planning (Ma et al. 2019), are also widespread.
Recently, learning-based MAPF solvers gained attention. One of the first such successful solvers was PRIMAL (Sartoretti et al. 2019) that demonstrated how MAPF problem can be solved in a decentralized manner utilizing machine learning. The recent learnable MAPF solvers such as SCRIMP (Wang et al. 2023), DCC (Ma, Luo, and Pan 2021), Follower (Skrynnik et al. 2024a), to name a few, typically rely on reinforcement learning and on additional modules, like the communication one, to solve the problem at hand. Orthogonally to these approaches, we rely purely on imitation learning from expert data.
@@
Offline deep reinforcement learning develops a policy based on previously collected data without interacting with the environment while training (Levine et al. 2020). This allows getting a robust policy due to the utilization of large amounts of pre-collected data. There are numerous effective offline RL approaches, such as CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2022), TD3+BC (Fujimoto and Gu 2021). Modern approaches often involve transformers as the architectural backbone. One popular approach is the Decision Transformer (DT) (Chen et al. 2021), which models the behavior of an expert by conditioning on the desired outcomes, thereby integrating reward guidance directly into the decision-making process. In multi-agent scenarios, there is less diversity in offline RL methods; however, a multi-agent adaptation of the DT exists, known as MADT (Meng et al. 2021).
Multi-agent imitation learning (MAIL)
@@
Imitation learning and learning from demonstration are actively used in multi-agent systems (Tang et al. 2024; Liu and Zhu 2024). MAIL refers to the problem of agents learning to perform a task in a multi-agent system through observing and imitating expert demonstrations without any knowledge of a reward function from the environment. It has gained particular popularity in the tasks of controlling urban traffic and traffic lights at intersections (Bhattacharyya et al. 2018; Huang et al. 2023) due to the presence of a large amount of data collected in real conditions and a high-quality simulator (such as Sumo (Lopez et al. 2018)). Among the methods in the field of MAIL, it is possible to note works using the Bayesian approach (Yang et al. 2020), generative adversarial methods (Song et al. 2018; Li et al. 2024), statistical tools for capturing multi-agent dependencies (Wang et al. 2021), low-rank subspaces (Shih, Ermon, and Sadigh 2022), latent multi-agent coordination models (Le et al. 2017), decision transformers (Meng et al. 2021), etc. Demonstrations are often used for pre-training in game tasks, such as training models for chess (Silver et al. 2016; Ruoss et al. 2024), and in MAPF tasks, as exemplified by SCRIMP (Wang et al. 2023). However, despite the listed works in this area, a single foundation model has not yet been proposed, the imitation learning, which already gives high results in multi-agent tasks and does not require an additional stage of online learning in the environment. This is largely due to the complexity of the behavioral multi-agent policies in various tasks (such as in StarCraft (Samvelyan et al. 2019) and traffic control) and the lack of large datasets of expert trajectories, which are necessary for effective training of foundation models. In this regard, the MAPF task is a convenient testbed for investigating transformer foundation models in a multi-agent setting, which will provide additional insights for creating such models in other applications, and our work also provides a large dataset for training MAPF models.
@@
The classical variant of the MAPF problem is defined by a tuple $(n,\mathcal{G=(V,E)},S=\{s_{1},...,s_{n}|~{}s_{i}\in\mathcal{V}\},G=\{g_{1},...,g_{n}|~{}g_{i}\in\mathcal{V}\})$, where $n$ is the number of agents acting in the shared workspace which is represented as an undirected graph $\mathcal{G}$. At each time step, an agent is assumed to either move from one vertex to the other or wait at the current vertex. The duration of both actions is uniform and equals $1$ time step. The plan for the $i$-th agent, $pl_{i}$, is a sequence of moves, s.t., each move starts where the previous one ends. Two distinct plans have a vertex (or edge) conflict if, at any time step, the agents occupy the same vertex (or traverse the same edge in opposite directions) at that time.
The task is to find a set of $n$ plans, $Pl=\{pl_{1},...,pl_{n}\}$, s.t. each $pl_{i}$ starts at $s_{i}$, ends at $g_{i}$ and each pair of plans in $Pl$ is conflict-free. The objective to be minimized is typically defined as $SoC(Pl)=\sum_{i=1}^{n}cost(pl_{i})$ (called the Sum-Of-Costs) or as $MS(Pl)=\max_{i=1,...,n}cost(pl_{i})$ (called the Makespan), where $cost(pl_{i})$ is the cost of the individual plan which equals the time step when the agent reaches its goal vertex (and does not move away further on).
@@
Despite MAPF being typically considered to be a planning problem as defined above, it can also be considered as a sequential decision-making (SDM) problem. Within the SDM framework, the problem is to construct a policy, $\pi$, that is a function that maps the current state (the current positions of all agents in the graph) to a (joint) action $\mathbf{a}=a_{i}\times...\times a_{n}$, where $a_{i}\in A_{i}$ and $A_{i}$ is the set of possible actions for agent $i$. When $\pi$ is obtained, it is invoked sequentially until either all agents reach their goals or the threshold on the number of time steps, $t_{max}$, is reached.
For better scalability, the decision-making policy might be decentralized, i.e., each agent chooses its action independently of the other agents. In practice, decentralized agents typically don’t have access to the global state of the environment, i.e., positions of the other agents, but rather rely on local observation, $o_{t}$. For example, if the underlying graph is a 4-connected grid, then the local observation may be a $(2r+1)\times(2r+1)$ patch of the grid centered at the agent’s current position (where $r$ is the observation radius) and the latter observes only the agents that are within this patch. A sequence of individual observations and actions forms the agent’s history: $h_{t}=\{o_{1},a_{1},o_{2},a_{2},...,o_{t-1},a_{t-1},o_{t}\}$, where $o_{k}$ and $a_{k}$ denote the action and the observation at time step $k$. This history is typically used to reconstruct a Markovian state of the environment via some approximator $f$: $s_{t}\approx f(h_{t})$ (e.g. $f$ can be represented as a neural network).
@@
To construct (learn) a decision making policy $\pi$, imitation learning relies on an expert policy $\pi^{b}$, which is used to collect a set of trajectories: $\mathcal{D}=\{traj\}$, where each trajectory is composed of the observations and actions: $traj=\{o^{b}_{1},a^{b}_{1},\dots,o^{b}_{L},a^{b}_{L}\}$. Intuitively, in MAPF context, $\mathcal{D}$ represents the expert knowledge on how an agent should behave under different circumstances (it can be obtained running a well-established MAPF solver on a range of problem instances).
@@
where $a_{j}\sim\pi_{\theta}(s_{j})$, $a^{b}_{j}\sim\pi_{\theta}(s^{b}_{j})$, and $\mathcal{L}$ is a loss function, which is selected depending on the action space. In the considered case, when the action space is discrete, cross-entropy is widely used as $\mathcal{L}$.
@@
Figure 1: The general pipeline of the MAPF-GPT: (1) Creating MAPF scenarios. (2) Generating ground truth data, i.e. MAPF solutions using an expert solver. (3) Transforming the solutions to the observation-action pairs and tokenization of the observations, which converts them into a format suitable for transformer architectures. (4) Executing the training loop, where observation/action pairs are sampled from the dataset, and the model is trained using cross-entropy loss.
@@
A large, curated dataset is crucial for any data-driven method including ours. To create the set of training instances we used POGEMA (Skrynnik et al. 2025), a versatile tool for developing learnable MAPF solvers that includes utilities to generate maze-like maps and maps with random obstacles, as well as to create MAPF instances from them (i.e. assigning start-goal locations). For our purposes we generate 10K of maze-like maps and 2.5K random maps and further created 3.75M different problem instances on these maps. The size of the maps varies from $17\times 17$ to $21\times 21$, the number of the agents is $16$, $24$, or $32$. Please note that as we aim to create an individual policy to solve MAPF in a decentralized fashion (i.e. each agent makes its own decision on how to move based on its local observation) it is not the size of the maps that actually matters but rather the density, i.e., the ratio of the free space to the space occupied by the agents. We use moderate and considerably high densities to make the agents face challenging patterns requiring coordination (especially on the maze-like maps).
@@
To create expert data, we use a recent variant of LaCAM (Okumura 2024, 2023), a state-of-the-art MAPF solver that is tailored to quickly find a solution and iteratively enhance it while having a time budget. As we need to solve a large number of MAPF instances (i.e. 3.75M) we set this time budget to be $10$ seconds.
The output of LaCAM on a single MAPF instance is the set of the individual plans. We further trace each plan and reconstruct an agent’s (local) observations in order to form observation-action pairs. We use the following post-processing to filter out some of them. First, if several pairs share the exactly the same observation we keep only one of them (picked randomly). Second, we observe that the fraction of the pairs when an agents waits at the goal is very high, because in many cases numerous agents wait for long times for other agents to reach their goals and just stand still. In fact, almost $40\%$ of the actions in the original expert data are the waiting ones. To remove this imbalance we discard $80\%$ of wait-at-target actions. We end up with 900M observation-action pairs from the maze-like maps, and 100M – from the random ones. A 9:1 proportion is chosen due to the maze maps possessing more challenging layouts with numerous narrow passages that require a high degree of cooperation between the agents.
We believe that the obtained dataset composed of 1B observation-action pairs is currently the largest dataset of such kind and may bring value to the other researchers developing learnable MAPF solvers. Additional technical details are presented in the Appendix C of the paper.
@@
Figure 2: The tokenization process for the MAPF-GPT model uses a vocabulary of 67 tokens, with an input of 256 tokens. Fewer tokens are shown for clarity and visibility.
Tokenization can be thought of as the process of transforming the data, observation-action pairs in our case, into the sequence of special symbols, tokens, to be further fed to the neural network that is trained to predict a single token, i.e. action, from the sequence of the input tokens, i.e. the ones that encode the observation. The input tokens are typically referred to as the context. We now wish to describe how our context, i.e. the observation, is structured.
@@
The second part of the observation contains data about the agent itself and the nearby ones. The information about each agent consists of the coordinates of its current and goal locations, actions history, i.e., the actions that were made in the previous $k$ steps, and an action that is preferable w.r.t. the agent’s individual cost-to-go map – the so-called greedy action. Please note that there may be cases where more than one action leads to a decrease in cost-to-go. Thus, we use special markers to indicate these multi-direction greedy actions (e.g., “up-right”).
@@
The rest of the input ($135$ tokens) is devoted to the information about agents. As it’s important to consider only the agents that can potentially influence the egocentric agent, we consider only the ones that are located in the $11\times 1$ 11 field of view at the current time step. The information about each agent is encoded via $10$ tokens: $2$ for the current position, $2$ for goal location, $5$ for action history, and $1$ for the next greedy action. Thus, we are able to encode the information about at most $13$ agents, including the egocentric one. The rest of $5$ tokens in the input are encoded with the empty token. In case there are not enough agents in the local field of view, the information about missing agents is also filled with empty tokens. The information about agents is sorted based on the distance to the egocentric agent, i.e., the information about the egocentric agent itself always goes first, as its distance is always 00.
@@
Figure 3: Success rate of the evaluated MAPF solvers on different maps. The shaded area indicates $95\%$ confidence intervals.
Figure 4: Quality of the obtained solutions relative to the ones of LaCAM (lower is better).
@@
As the model backbone, we used a modern decoder-only transformer (Brown et al. 2020). We used a softmax layer to parameterize the discrete probability distribution. To sample an action, we then used multinomial sampling. The length of the input sequence (context size) of the model is $256$. The output size of the model is $5$ since each agent has $5$ available discrete actions. We used learnable position embeddings. We don’t use causal masking, which is common practice in the NLP (Radford et al. 2019), since the model predicts only a single action ahead in a non-autoregressive manner. To speed up training, we used the flash attention technique (Dao 2024).
We use models of different sizes (i.e. number of parameters) in our work. Specifically, our largest model contains $85$M parameters. We also consider much smaller models that contain $6$M and $2$M parameters.